OpenStudy (anonymous):

Sue used a probability simulator to roll a 6-sided number cube and flip a coin 100 times. The results are shown in the tables below: Number on the Cube Number of Times Rolled 1 18 2 25 3 12 4 32 5 3 6 10 Heads Tails 56 44 Using Sue's simulation, what is the probability of rolling a 2 on the number cube and the coin landing heads up?

OpenStudy (anonymous):

Answers are 81/100 56/100 1400/100 and 1100/100

OpenStudy (anonymous):

MEDAL TO ANYONE WHO HELPS ME

OpenStudy (anonymous):

do \[\frac{ 25 }{ 100 }*\frac{ 56 }{ 100 }\]

OpenStudy (anonymous):

then simplify your answer

OpenStudy (anonymous):

1400

OpenStudy (anonymous):

1400/100

OpenStudy (anonymous):

no, 1400/10000 is correct

OpenStudy (anonymous):

wait simplify what

OpenStudy (anonymous):

14

OpenStudy (anonymous):

can you take a screenshot of the options

OpenStudy (anonymous):

don't simplify yet

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

oen sec

OpenStudy (anonymous):

is \[\frac{ 1400 }{ 10,000 }\] one of your options?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, then that is your answer

OpenStudy (anonymous):

oh ok wel,l thank you